OpenStudy (anonymous):

1 1/7* 9 3/4= & simplify the answer using whole numbers, proper fraction, or mixed number.

jhonyy9 (jhonyy9):

8/7 * 39/4 = (8*39)/(7*4) =312/28

OpenStudy (anonymous):

Thank you

jhonyy9 (jhonyy9):

= 156/14 =78/7

jhonyy9 (jhonyy9):

yw

jhonyy9 (jhonyy9):

good luck

jhonyy9 (jhonyy9):

bye

OpenStudy (anonymous):

which answer is correct 3 12/28 or 78/7

OpenStudy (anonymous):

\[\left(1\frac{1}{7}\right)*\left(9\frac{3}{4}\right)=\frac{78}{7}=11\frac{1}{7} \]

OpenStudy (anonymous):

Thanks